# MD-NOMAD: Mixture density nonlinear manifold decoder for emulating stochastic differential equations and uncertainty propagation

### v1

@@ 5.1 Stochastic Van der Pol equation
$\text{where}\;X_{0}$ $=-3,\;Y_{0}=0,\;\lambda\sim\mathcal{U}(0.4,0.8),\;t\in[0,20].$
@@
We use the MD-NOMAD to learn the mapping from $[\lambda,t]$ to $[\pi(\lambda)(t),\,\mu(\lambda)(t),\,\sigma(\lambda)(t)]$. The mapping is eventually employed to obtain the probability density function $f_{X|\Lambda,T}(x)(\lambda,t)$. We use a small two-layered FCN for the branch network and a comparatively larger 4-layered FCN for the decoder network. In addition, the stochastic noise strength is employed as an input to the branch net whereas $t$ is used as the augmenting input to the decoder network. Due to our choice for the experimental design, the complete dataset consists of $\mathscr{N}\times\mathscr{R}$ solution trajectories. Furthermore, the full dataset is randomly split into training and validation datasets with a $9:1$ size ratio. The Adam optimizer is employed for the mini-batch-based training with a suitable learning rate and schedule, while the number of training epochs and mixture components is set to $300$ and $15$ respectively. Also, the number of layers in the branch and decoder network along with the optimizer and the train-validation split is kept the same, unless stated otherwise, for different problems assessed in this study. In addition, the number of realizations and replications used for the current example are $\mathscr{N}=50$ and $\mathscr{R}=40$, respectively.
@@
Finally, to assess the performance of the trained MD-NOMAD, we perform a comparison between the reference distribution and statistics with the model-predicted PDF and the statistics. We carry out this comparison on two randomly chosen values for stochastic noise strength $\lambda=0.45$ and $\lambda=0.75$, which are not present in the training or the validation dataset. Specifically, two types of pictorial comparisons are performed — the first one being the comparison between the predicted PDF and the reference distribution for $X_{t}$ at a few different temporal locations through the solution time span to ascertain how well the proposed model performs in capturing multiple moments of the desired distribution, and the second one is the comparison between the reference statistics, i.e., mean and standard deviation, and the predicted statistics for the entire solution time-span. Both the reference distribution and corresponding statistics are obtained from the empirical distribution of $10^{4}$ replications. We again emphasize that as the output of MD-NOMAD is the parameters for a mixture of Gaussian distribution, the predicted PDF and statistics are obtained in an analytic way obviating the need for a large number of MC simulation (MCS) runs. The results for comparison between the predicted PDF and the reference distribution at different temporal locations are presented in Figure 2, while the comparison between the temporal evolution of predicted and reference statistics is presented in Figure 3. It can be observed from Figure 2 that MD-NOMAD can predict PDFs for $X_{t}$ with reasonable accuracy even after the bifurcation, i.e., where the PDF transitions from unimodal to bimodal.
@@
5.2 N-dimensional stochastic Lorenz-96
The stochastic Lorenz-96 system represents a nonlinear set of stochastic ordinary differential equations, offering a simplified model to capture atmospheric dynamics, particularly the intricate multi-scale interactions. We consider the $N$-dimensional stochastic Lorenz-96 system, expressed as:
@@ 5.2 N-dimensional stochastic Lorenz-96
where $i\in\{1,2,\ldots,N\}$ denotes cyclic indices. For this illustration, we aim to construct a surrogate for all $N$-components of the state variable $X_{t}$. The trajectories for the Lorenz-96 system are generated utilizing the Euler-Maruyama method with a time step $dt=0.04$ over the period $t=[0,4]s$. In addition, the experimental design is based on the diffusion coefficient $\lambda$, which is sampled from a uniform distribution $\mathcal{U}(0.15,0.35)$ with $\mathscr{N}=50$ realizations and $\mathscr{R}=30$ replications. Despite the introduction of a general $N$-dimensional system, our investigation confines itself to two scenarios: $N=10$ and $N=50$. It’s important to recognize that this system also experiences bifurcation, and the PDFs for different components of $X_{t}$ transition from densities with sharper peaks to diffused ones as time progresses.
@@
The MD-NOMAD, for this system of equations, is used to construct a surrogate for the following mapping: $[\lambda,t,x_{d}]$ $\mapsto$ $[\pi(\lambda)(t,x_{c}),\,\mu(\lambda)(t,x_{c}),\,\sigma(\lambda)(t,x_{c})]$. We use $\lambda$ as the branch network’s input, while $(t,x_{d})$ are concatenated to latent branch network outputs to form the inputs for the decoder network. As we employ a pointwise NN architecture in this study, we require conditioning variables for different components $X^{i}$ of the $N$-dimensional Lorenz equation. In this example, $x_{d}\in[1,2,\ldots,N]$ acts as a conditioning variable for enabling MD-NOMAD to construct the required surrogate map for a particular component. It is important to understand that $x_{d}$ has to be scaled suitably before supplying it to the decoder network. We scale $x_{d}$ by dividing the size of Lorenz’s system, i.e., $N$. Furthermore, during the preprocessing of the training and validation dataset, the system dynamics trajectories are temporally subsampled by a factor of $2$ to mimic a setting with sparser data availability. The number of components used for the Gaussian mixture and the total number of training is chosen to be $15$ and $250$ for the $50-$dimensional system, whereas, for the $10-$dimensional system these are set to $10$ and $250$. For conciseness, the results containing the predictions for the PDFs at different times, which can be found in Figure 4 and Figure 5, are presented for some selected components. Furthermore, the results containing the statistics are presented in Figure 6 and Figure 7. The results correspond to a case with stochastic noise strength $\lambda=0.22$, which is not a part of the training and validation dataset. Furthermore, the reference distribution and corresponding statistics are obtained from the empirical distribution of $10^{4}$ replications.
@@
Figure 4: Comparison between reference distribution and MD-NOMAD predicted PDF at times $t={0.22,\,1.76,\,2.96,\,\text{and }4.0}$ s corresponding to $\lambda=0.22$ for the (a) second, (b) fifth, and (c) tenth solution component of the $10$-d stochastic Lorenz equations.
@@
Figure 5: Comparison between reference distribution and MD-NOMAD predicted PDF at times $t={0.22,\,1.76,\,2.96,\,\text{and }4.0}$ s corresponding to $\lambda=0.22$ for the (a) second, (b) twenty fifth, and (c) fiftieth solution component of the $50$-d stochastic Lorenz equations.
@@
Figure 6: Comparison between reference statistics and MD-NOMAD predicted statistics till $t=4$ s for all ten components ($X^{i}$) of the $10$-d stochastic Lorenz equations corresponding to the test case with $\lambda=0.22$.
@@
Figure 7: Comparison between reference statistics and MD-NOMAD predicted statistics till $t=4$ s for all ten components ($X^{i}$) of the $50$-d stochastic Lorenz equations corresponding to the test case with $\lambda=0.22$. For compendiousness, the mean and standard deviation values, for a time span of 4 s and all the components (in an ascending component number order), are vertically stacked to obtain a 2-d image.
@@ 5.3 Bimodal analytical benchmark
$f_{Y|X,\Lambda}(y)(X=x,\Lambda=\lambda)=\lambda~{}\varphi(4\sin^{2}(\pi\cdot x)+4x-2)+(1-\lambda)~{}\varphi(4\sin^{2}(\pi\cdot x)-4x+2).$ (11)
Here, $\varphi$ denotes a normal PDF with a nonlinear mean function with respect to $x$ and a standard deviation of $0.8$. The parameter $\lambda$ parameterizes the mixture coefficients, and $X$ is an input vector of $x$-values consisting of $100$ uniformly spaced points between 00 and $1$. The target PDF is a mixture of two Gaussian PDFs. The experimental design is based on the mixture coefficient parameter $\lambda$, which is sampled from a uniform distribution $\mathcal{U}(0.4,0.7)$ with $\mathscr{N}=70$ realizations and $\mathscr{R}=30$ replications. The number of training epochs and the number of mixture component is equal to $300$ and $10$ respectively. The objective is to construct the following mapping with MD-NOMAD: $[\lambda,x]$ $\mapsto$ $[\pi(\lambda)(x),\,\mu(\lambda)(x),\,\sigma(\lambda)(x)]$. The complementary pair $(\lambda,1-\lambda)$ is used as an input to the branch network, while $x$ is appended to the latent output of the branch network to create an input for the decoder network.
@@
Model $\ell_{xy}=(0.3,0.15)$ $\ell_{xy}=(0.06,0.12)$
@@
5.5 1-dimensional unsteady stochastic heat equation
@@
5.6 1-dimensional unsteady stochastic viscous Burgers equation
@@ 6 Conclusions
We introduced the mixture density-based NOMAD framework, which seamlessly integrates pointwise operator learning network NOMAD with mixture density formalism, enabling the estimation of conditional probability distributions for stochastic output functions. The components of the mixture were Gaussian distributions, which were predominantly employed for their mathematical traceability. The scalability inherited from the NOMAD architecture to handle high-dimensional problems, coupled with the ability to learn a data representation of stochastic systems, underscores the versatility and practical utility of MD-NOMAD. In our empirical assessment, we validated the efficacy of MD-NOMAD across a diverse range of stochastic ordinary and partial differential equations (including high-dimensional examples), covering both linear and nonlinear formulations, alongside analytical benchmark problems. Furthermore, the usage of MD-NOMAD was also demonstrated for one-shot uncertainty propagation in a 2-dimensional stochastic partial differential equation. In addition, as the predicted PDFs and associated statistics for MD-NOMAD can be computed analytically, without resorting to MC simulations, the proposed framework experiences an improved computational efficiency, thereby bolstering its practical applicability. However, it is important to note that MD-NOMAD necessitates additional attention to adjusting an extra hyperparameter compared to deterministic NOMAD, specifically, the number of mixture components. Presently, the determination of the number of mixture components relies on validation error. However, Bayesian decision-making can be leveraged for automatic selection of the number of mixture components in later studies. Furthermore, future investigations into alternative architectures beyond the FCN utilized in this study, along with further endeavors such as assessing the capabilities of MD-NOMAD in constructing surrogates for complex 3-dimensional problems like turbulent channel flow, would be of considerable interest.
7 Acknowledgement

### v2

@@ 5.1 Stochastic Van der Pol equation
$\text{where}\;X_{0}$ $=-3,\;Y_{0}=0,\;\mu=1,\;\lambda\sim\mathcal{U}(0.4,0.8),\;t\in[0,20].$
@@
We use the MD-NOMAD to learn the mapping from $[\lambda,t]$ to $[\pi(\lambda)(t),\,\mu(\lambda)(t),\,\sigma(\lambda)(t)]$. The mapping is eventually employed to obtain the probability density function $f_{X|\Lambda,T}(x)(\lambda,t)$. We use a small two-layered fully connected network (FCN) for the branch network and a comparatively larger 4-layered FCN for the decoder network. In addition, the stochastic noise strength is employed as an input to the branch net whereas $t$ is used as the augmenting input to the decoder network. Due to our choice for the experimental design, the complete dataset consists of $\mathscr{N}\times\mathscr{R}$ solution trajectories. Furthermore, the full dataset is randomly split into training and validation datasets with a $9:1$ size ratio. The Adam optimizer is employed for the mini-batch-based training with a suitable learning rate and schedule, while the number of training epochs and mixture components is set to $300$ and $15$ respectively. Also, the number of layers in the branch and decoder network along with the optimizer and the train-validation split is kept the same, unless stated otherwise, for different problems assessed in this study. In addition, the number of realizations and replications used for the current example are $\mathscr{N}=50$ and $\mathscr{R}=40$, respectively.
@@
Finally, to assess the performance of the trained MD-NOMAD, we perform a comparison between the reference distribution and statistics with the model-predicted PDF and the statistics. We carry out this comparison on two randomly chosen values for stochastic noise strength $\lambda=0.45$ and $\lambda=0.75$, which are not present in the training or the validation dataset. Specifically, two types of pictorial comparisons are performed — the first one being the comparison between the predicted PDF and the reference distribution for $X_{t}$ at a few different temporal locations through the solution time span to ascertain how well the proposed model performs in capturing multiple moments of the desired distribution, and the second one is the comparison between the reference statistics, i.e., mean and standard deviation, and the predicted statistics for the entire solution time-span. Both the reference distribution and corresponding statistics are obtained from the empirical distribution of $10^{4}$ replications. The samples are generated using the same stochastic simulator utilized for producing the training data, both for the current case and for all other example problems presented in this study. We again emphasize that as the output of MD-NOMAD is the parameters for a mixture of Gaussian distribution, the predicted PDF and statistics are obtained in an analytic way obviating the need for a large number of MC simulation (MCS) runs. The results for comparison between the predicted PDF and the reference distribution at different temporal locations are presented in Figure 2, while the comparison between the temporal evolution of predicted and reference statistics is presented in Figure 3. It can be observed from Figure 2 that MD-NOMAD can predict PDFs for $X_{t}$ with reasonable accuracy even after the bifurcation, i.e., where the PDF transitions from unimodal to bimodal.
@@
5.2 N-component stochastic Lorenz-96
The stochastic Lorenz-96 system represents a nonlinear set of stochastic ordinary differential equations, offering a simplified model to capture atmospheric dynamics, particularly the intricate multi-scale interactions. We consider the $N$-component stochastic Lorenz-96 system, expressed as:
@@ 5.2 N-component stochastic Lorenz-96
where $i\in\{1,2,\ldots,N\}$ denotes cyclic indices. For this illustration, we aim to construct a surrogate for all $N$-components of the state variable $X_{t}$. The trajectories for the Lorenz-96 system are generated utilizing the Euler-Maruyama method with a time step $dt=0.04$ over the period $t=[0,4]s$. In addition, the experimental design is based on the diffusion coefficient $\lambda$, which is sampled from a uniform distribution $\mathcal{U}(0.15,0.35)$ with $\mathscr{N}=50$ realizations and $\mathscr{R}=30$ replications. Despite the introduction of a general $N$-component system, our investigation confines itself to two scenarios: $N=10$ and $N=50$. It’s important to recognize that this system also experiences bifurcation, and the PDFs for different components of $X_{t}$ transition from densities with sharper peaks to diffused ones as time progresses.
@@
The MD-NOMAD, for this system of equations, is used to construct a surrogate for the following mapping: $[\lambda,t,x_{d}]$ $\mapsto$ $[\pi(\lambda)(t,x_{c}),\,\mu(\lambda)(t,x_{c}),\,\sigma(\lambda)(t,x_{c})]$. We use $\lambda$ as the branch network’s input, while $(t,x_{d})$ are concatenated to latent branch network outputs to form the inputs for the decoder network. As we employ a pointwise NN architecture in this study, we require conditioning variables for different components $X^{i}$ of the $N$-component Lorenz equation. In this example, $x_{d}\in[1,2,\ldots,N]$ acts as a conditioning variable for enabling MD-NOMAD to construct the required surrogate map for a particular component. It is important to understand that $x_{d}$ has to be scaled suitably before supplying it to the decoder network. We scale $x_{d}$ by dividing the size of Lorenz’s system, i.e., $N$. Furthermore, during the preprocessing of the training and validation dataset, the system dynamics trajectories are temporally subsampled by a factor of $2$ to mimic a setting with sparser data availability. The number of components used for the Gaussian mixture and the total number of training epochs is chosen to be $15$ and $250$ for the $50-$component system, whereas, for the $10-$component system these are set to $10$ and $250$. For conciseness, the results containing the predictions for the PDFs at different times, which can be found in Figure 4 and Figure 5, are presented for some selected components. Furthermore, the results containing the statistics are presented in Figure 6 and Figure 7. The results correspond to a case with stochastic noise strength $\lambda=0.22$, which is not a part of the training and validation dataset. Furthermore, the reference distribution and corresponding statistics are obtained from the empirical distribution of $10^{4}$ replications.
@@
Figure 4: Comparison between reference distribution and MD-NOMAD predicted PDF at times $t={0.22,\,1.76,\,2.96,\,\text{and }4.0}$ s corresponding to $\lambda=0.22$ for the (a) second, (b) fifth, and (c) tenth solution component of the $10$-component stochastic Lorenz equations.
@@
Figure 5: Comparison between reference distribution and MD-NOMAD predicted PDF at times $t={0.22,\,1.76,\,2.96,\,\text{and }4.0}$ s corresponding to $\lambda=0.22$ for the (a) second, (b) twenty fifth, and (c) fiftieth solution component of the $50$-component stochastic Lorenz equations.
@@
Figure 6: Comparison between reference statistics and MD-NOMAD predicted statistics till $t=4$ s for all ten components ($X^{i}$) of the $10$-component stochastic Lorenz equations corresponding to the test case with $\lambda=0.22$.
@@
Figure 7: Comparison between reference statistics and MD-NOMAD predicted statistics till $t=4$ s for all ten components ($X^{i}$) of the $50$-component stochastic Lorenz equations corresponding to the test case with $\lambda=0.22$. For compendiousness, the mean and standard deviation values, for a time span of 4 s and all the components (in an ascending component number order), are vertically stacked to obtain a 2-d image.
@@ 5.3 Bimodal analytical benchmark
$f_{Y|X,\Lambda}(y)(X=x,\Lambda=\lambda)=\lambda~{}\mathit{\phi}(4\sin^{2}(\pi\cdot x)+4x-2)+(1-\lambda)~{}\mathit{\phi}(4\sin^{2}(\pi\cdot x)-4x+2).$ (17)
Here, $\mathit{\phi}$ denotes a normal PDF with a nonlinear mean function with respect to $x$ and a standard deviation of $0.8$. The parameter $\lambda$ parameterizes the mixture coefficients, and $X$ is an input vector of $x$-values consisting of $100$ uniformly spaced points between 00 and $1$. The target PDF is a mixture of two Gaussian PDFs. The experimental design is based on the mixture coefficient parameter $\lambda$, which is sampled from a uniform distribution $\mathcal{U}(0.4,0.7)$ with $\mathscr{N}=70$ realizations and $\mathscr{R}=30$ replications. The number of training epochs and the number of mixture component is equal to $300$ and $10$ respectively. The objective is to construct the following mapping with MD-NOMAD: $[\lambda,x]$ $\mapsto$ $[\pi(\lambda)(x),\,\mu(\lambda)(x),\,\sigma(\lambda)(x)]$. The complementary pair $(\lambda,1-\lambda)$ is used as an input to the branch network, while $x$ is appended to the latent output of the branch network to create an input for the decoder network.
@@
Model $\boldsymbol{\ell_{xy}=\left(0.3,0.15\right)}$ $\boldsymbol{\ell_{xy}=\left(0.06,0.12\right)}$
@@
5.4.1 Computational cost comparison between the MC method, KDE, and MD-NOMAD
In this section, we evaluate and compare the computational cost associated with the Monte Carlo (MC) method, Kernel Density Estimation (KDE), and MD-NOMAD for estimating probability density functions (PDFs) and statistical quantities, with computational time serving as the primary metric. In general, statistical computations using the MC method require a large number of samples, typically on the order of $10^{4}$. Consequently, to estimate statistics at a given spatial location $(x,y)$ for the two-dimensional SPDE, substantial sample generation is necessary. Similarly, KDE also relies on a large sample size to produce accurate and smooth PDF estimates. In contrast, MD-NOMAD eliminates the need for sample generation altogether, thereby achieving a significant reduction in computational overhead. This advantage is clearly reflected in the wall-clock times reported for each method in the estimation of PDFs and statistical measures, as presented in Table 2.
5.4.2 Demonstration of resolution invariance in MD-NOMAD
In this section, we demonstrate the resolution invariance of MD-NOMAD by evaluating its performance on a different grid resolution from that used during training. Specifically, the model was originally trained on a $32\times 32$ grid. We now test it on an UP problem defined over a $15\times 15$ grid, with spatial locations that do not overlap with the training grid. Figure 12 presents the predicted mean fields for both the training and testing grids corresponding to the length scale $\ell_{xy}=(0.3,0.15)$, along with the grids.
Figure 12: Illustration of resolution invariance in MD-NOMAD. (Left) Comparison of the training and testing grids: black crosses denote the original $32\times 32$ training grid points, and red circles indicate the $15\times 15$ testing grid points with different spatial locations. (Middle) Predicted mean field on the $15\times 15$ testing grid. (Right) Predicted mean field on the original $32\times 32$ training grid. Mean fields correspond to the length scale $\ell_{xy}=(0.3,0.15)$.
Table 2: Computational time comparison between MC-method, KDE, and MD-NOMAD for PDF and statistical computations for the 2-dimensional SPDE.
Method PDF (s) Statistics (s)
MD-NOMAD 0.037 0.398
KDE and MC 301.406 301.847
5.5 1-dimensional stochastic heat equation
@@ 5.4.2 Demonstration of resolution invariance in MD-NOMAD
5.6 1-dimensional stochastic viscous Burgers equation
@@
Table 3: Comparison of expected Wasserstein distance of second-order ($\mathcal{E}_{W}$) and KL divergence ($\mathcal{E}_{\text{KL}}$) for MD-NOMAD and KCDE-based model across all example problems on an unseen test set.
Problem MD-NOMAD KCDE
$\mathcal{E}_{W}$ $\mathcal{E}_{\text{KL}}$ $\mathcal{E}_{W}$ $\mathcal{E}_{\text{KL}}$
Stochastic Van der Pol $2.14\times 10^{-1}$ $7.33\times 10^{-2}$ $6.09\times 10^{-1}$ $2.81\times 10^{-1}$
10-component stochastic Lorenz-96 $7.26\times 10^{-4}$ $2.64\times 10^{-3}$ $4.10\times 10^{-3}$ $1.05\times 10^{-1}$
50-component stochastic Lorenz-96 $6.28\times 10^{-4}$ $2.19\times 10^{-3}$ $3.76\times 10^{-3}$ $8.18\times 10^{-2}$
Bimodal analytical benchmark $1.55\times 10^{-1}$ $1.89\times 10^{-3}$ $4.58\times 10^{-1}$ $2.74\times 10^{-2}$
2-d SPDE $1.02\times 10^{-4}$ $8.12\times 10^{-3}$ $1.24\times 10^{-1}$ $2.34$
1-d stochastic heat $5.46\times 10^{-5}$ $6.05\times 10^{-3}$ $3.24\times 10^{-4}$ $5.45\times 10^{-2}$
1-d stochastic viscous Burgers $1.60\times 10^{-5}$ $4.23\times 10^{-3}$ $8.88\times 10^{-3}$ $2.02$
@@ 6 Conclusions
We introduced the mixture density-based NOMAD framework, which seamlessly integrates pointwise operator learning network NOMAD with mixture density formalism, enabling the estimation of conditional probability distributions for stochastic output functions. The components of the mixture were Gaussian distributions, which were predominantly employed for their mathematical traceability. The scalability inherited from the NOMAD architecture to handle high-dimensional problems, coupled with the ability to learn a data representation of stochastic systems, underscores the versatility and practical utility of MD-NOMAD. In our empirical assessment, we validated the efficacy of MD-NOMAD across a diverse range of stochastic ordinary and partial differential equations (including high-dimensional examples), covering both linear and nonlinear formulations, alongside analytical benchmark problems. In addition, MD-NOMAD was evaluated against the KCDE-based emulator, and as demonstrated in Table 3, MD-NOMAD consistently outperforms KCDE across all benchmark problems. It is important to note that training a KCDE-based model becomes computationally demanding and increasingly complex for high-dimensional problems because of the curse of dimensionality. Furthermore, the usage of MD-NOMAD was also demonstrated for one-shot uncertainty propagation in a 2-dimensional stochastic partial differential equation. In addition, as the predicted PDFs and associated statistics for MD-NOMAD can be computed analytically, without resorting to MC simulations, the proposed framework experiences an improved computational efficiency, thereby bolstering its practical applicability. However, it is important to note that MD-NOMAD necessitates additional attention to adjusting an extra hyperparameter compared to deterministic NOMAD, specifically, the number of mixture components. Presently, the determination of the number of mixture components relies on validation error. However, Bayesian decision-making can be leveraged for automatic selection of the number of mixture components in later studies. Also, in this article, we employed a mixture of univariate Gaussian. While this choice offers several practical advantages, we acknowledge that it may limit the ability to accurately capture inter-component covariances. Investigating this limitation could be a subject of future work. Furthermore, future investigations into alternative architectures beyond the FCN utilized in this study, along with further endeavors such as assessing the capabilities of MD-NOMAD in constructing surrogates for complex 3-dimensional problems like turbulent channel flow, would be of considerable interest. Moreover, extending MD-NOMAD to function as a surrogate model in practical reliability analysis tasks — such as for wind turbines, heritage structures, and composite plates — could also be an interesting direction for future research.
Acknowledgement